ACCESS(1) ACCESS(1) NAME access−determinewhetherafilecanbeaccessed SYNOPSIS access-modefile DESCRIPTION Exitsuccessfullyiffilecanbeaccessedwiththespecifiedmode.modeisoneormorelettersof rwx,whererisforreadable,wisforwritable,andxisforexecutable. Thedifferencebetweenaccessandtestisthatthelatterlooksatthepermissionbits,whilethe formerchecksusingtheaccess(2)systemcall.Thismakesadifferencewhenfilesystemshave beenmountedread-only. OPTIONS accessacceptsthefollowingadditionaloptions: --helpPrinthelpmessageandexit. --version Printversioninformationandexit. SEEALSO access(2) Kpathsea7.32x 4January1998 1